OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 2x2 - 2z4 + y2 - x2 + z4 if x = -4, y = 3, and z = 2. Numerical Answers Expected!

OpenStudy (anonymous):

#2 Evaluate the following expression using the values given: Find -a2 - 3b3 + c2 + 2b3 - c2 if a = 3, b = 2, and c = -3. -67 -17 -31 1

OpenStudy (lyrae):

Replace the letters with what they are equal to and do the aritmetic. So if i insert y = 3 i get \[2x^{2} - 2z ^{4} + 3^{2} - x ^{2} + z^{4}\] For negative numbers you insert them wrapped in a parentheses. For example \[2x ^{3}, x = -1\] becommes \[2 \times (-1)^{3} = -2\]